# Properties of Carbon-rich AGB Stars in the LMC and the Milky Way

### v2

@@ Abstract
We present a comparative study of carbon-rich asymptotic giant branch (CAGB) stars in the Large Magellanic Cloud (LMC; 7347 stars) and the Milky Way (7163 stars) using infrared color–magnitude diagrams (CMDs), spectral energy distributions (SEDs), two-color diagrams (2CDs), and variability data. Observed SEDs are compared with theoretical models to characterize the central stars and their circumstellar dust envelopes and to estimate distances. For the LMC, a set of best-fitting CAGB models is derived by fitting observed SEDs with radiative transfer models, utilizing the galaxy’s well-established distance. For Galactic CAGB stars, where Gaia DR3 parallaxes are uncertain, we estimate distances by fitting observed SEDs with the CAGB models validated against LMC stars, and for Mira variables, from the period–magnitude relation calibrated with LMC Miras. A comparison of these approaches demonstrates that the SED-based distances are both reliable and practical for a large sample of Galactic CAGB stars. We find that CAGB stars in both galaxies show broadly similar infrared properties, although the LMC sample lacks stars with extremely thick dust envelopes.
@@ 1 Introduction
LMC-CAGB_SAGE-Sb SJc - 151
LMC-CAGB_KDMK Kontizas et al. (2001) 7760d 24+15e
LMC-CAGB combined - 7347
@@
MW-CAGB combined - 7163
aSee Section 2.1. bSubgroup of LMC-CAGB_SAGE; objects identified from the Spitzer IRS. cSloan et al. (2016); Jones et al. (2017). d3933 objects are duplicates of LMC-CAGB_SAGE objects. A large fraction of this sample could be extrinsic carbon stars. eMiras from OGLE-III and candidates for new Miras identified from WISE data (Suh 2025).
@@
We analyze the properties of CAGB stars in the LMC and the Milky Way. Section 2 presents catalogs of known CAGB stars in both galaxies. Section 3 describes radiative transfer models for circumstellar dust shells and the derivation of model parameters, constrained by LMC CAGB stars. In Section 4, observed spectral energy distributions (SEDs) are compared with model predictions to characterize the central stars and their dust envelopes and to estimate theoretical distances. Section 5 compares observed infrared two-color diagrams (2CDs) and color–magnitude diagrams (CMDs) with theoretical models. Section 6 examines the period–magnitude relation (PMR) of Mira-type CAGB stars. Section 7 summarizes the main results.
2 Sample Stars
@@ 2.1 CAGB stars in the LMC
2.2 CAGB stars in the Milky Way
Suh (2024) compiled a catalog of CAGB stars in the Milky Way, classified into two groups: those with IRAS counterparts (MW-CAGB_IC), identified from the IRAS source catalog and generally corresponding to brighter or more isolated stars, and those without IRAS counterparts (MW-CAGB_NI), identified from the AllWISE or Gaia DR3 catalogs and typically representing fainter stars or objects in crowded regions. Most MW-CAGB_IC sources also have entries in the AllWISE or Gaia DR3 catalogs. In this study, we adopt the catalog of Galactic CAGB stars from Suh (2024) (MW-CAGB objects; see Table 1). The MW-CAGB_IC-S subset denotes those MW-CAGB_IC objects for which infrared spectral data are available (see Section 4).
@@ 2.2 CAGB stars in the Milky Way
2.3 Observational Data
Table 2 presents the visual and IR bands used in this study for constructing SEDs, CMDs and 2CDs. The table includes the reference wavelength ($\lambda_{ref}$) and zero-magnitude flux (ZMF) for each band, which are essential for deriving theoretical models (see Section 3) and building SEDs from observational data. For a comprehensive discussion of IR photometric data and ZMF values, refer to Section 2.1 of Suh (2021).
@@ 2.3 Observational Data
2.4 Distances and extinction for Galactic CAGB stars
@@
Gaia detectors operating in the visual bands cannot reliably detect CAGB stars with thick dust envelopes, often resulting in missing data for these objects. In addition, the Gaia counterparts provided by SIMBAD may be inaccurate for some objects, as it is often difficult to identify reliable counterparts using available photometric data and theoretical models, especially in crowded regions. Furthermore, CAGB stars frequently exhibit non-symmetric motions in their outer envelopes (e.g., Kim & Taam 2012), which can lead to inaccurate distance measurements in Gaia DR3. Therefore, distances derived from comparative analyses between observed and theoretical model SEDs may provide alternative or more reliable estimates (see Section 4).
@@
3 Theoretical Dust Shell Models
To calculate theoretical model SEDs for CAGB stars, we use radiative transfer models for spherically symmetric dust shells surrounding central stars. The radiative transfer code RADMC-3D (http://www.ita.uni-heidelberg.de/~dullemond/software/radmc-3d/), following the approaches of Suh & Kwon (2013) and Suh (2024), is applied. Detailed descriptions of the theoretical models and their limitations can be found in Suh (2020).
@@ 3 Theoretical Dust Shell Models
aSee Section 3 for details. For all models, $T_{c}$ = 1000 K. bSeven representative CAGB models were used to produce CMDs and 2CDs. cAMC: pure AMC; AMC+SIC: a simple mixture of AMC and SiC (20% by mass) dStellar luminosity in $10^{3}L_{\odot}$. Based on this luminosity, the scaled logarithmic distances can be sampled using a Gaussian Monte Carlo spread (GS) in $\log_{10}d$ ($\sigma=0.0798$ dex) derived from LMC-CAGB stars (see Figure 1), corresponding to a total luminosity variation of approximately $2.1\times$ from minimum to maximum (see Section 3.1).
3.1 Modeling LMC-CAGB stars
@@
For the full sample of 7347 LMC-CAGB stars (Table 1), we tested an extensive grid of radiative transfer models covering a broad range of parameter combinations to identify the most suitable set of CAGB models. The observed SEDs were fitted using a least-squares optimization method (see Sections 4 and 4.1).
Figure 1: The top panel shows CAGB model SEDs for seven representative models (see Table 3). For LMC-CAGB stars, the shows the distance distribution (middle) and the relations between dust optical depth and luminosity (bottom) derived from CAGB model SEDs (Table 3). See Section 3.1 for details.
Figure 2: For LMC-CAGB stars, the top panels show the relations between optical depth and luminosity derived using two different methods based on CAGB models. In the top panels, the areas of the filled circular markers are proportional to the square root of the number of objects at each point. The bottom panels present IR CMDs for LMC-CAGB stars, with the brown arrow indicating an object likely to be an extrinsic carbon star (see Section 4.1). For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$ = 0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 3.1 for details.
@@
The best-fitting CAGB model (from Table 3) and the corresponding scaling distance were derived for each LMC-CAGB star. Reliable results were obtained for 7205 of the 7347 LMC-CAGB stars (see Section 4). The bottom panel of Figure 1 shows the logarithmic distribution of the derived distances for the 7205 LMC-CAGB stars. The median distance (49.6 kpc) agrees well with the established value for the LMC, and the small dispersion in $\log_{10}d$ (0.0798 dex) demonstrates the consistency and reliability of the CAGB models (Table 3).
@@
The dispersion in $\log_{10}d$ (0.0798 dex; corresponding to roughly 20 % in distance; right panel of Figure 1) for LMC-CAGB stars likely arises from three factors: (1) positional variations within the LMC, whose radius is about 10 % of its mean distance, implying a $\sim$5 % dispersion in distance; (2) the fact that a given $\tau_{10}$ value can correspond to multiple luminosities, as stars with different initial masses may exhibit different luminosities; and (3) intrinsic luminosity variations caused by the pulsations of CAGB stars. The observed dispersion in $\log_{10}d$ can therefore be regarded as representing the overall uncertainty in luminosity, since the contribution of the first factor is relatively minor. The measured dispersion in $\log_{10}d$ corresponds to a luminosity dispersion of 0.160 dex for the CAGB models (see Table 3).
Figure 3: The top panels show IR CMDs for Galactic CAGB stars. For Galactic CAGB stars, distances were derived from CAGB model SEDs, with Gaussian Monte Carlo scatter (GS) in $\log_{10}d$ ($\sigma$$\pm$0.0798 dex; obtained from LMC-CAGB stars; see Figure 1). For CAGB stars in the LMC and the Milky Way, the middle panels show the relations between optical depth and luminosity, and the bottom panels show the luminosity functions. In the middle panels, the areas of the filled circular markers are proportional to the square root of the number of objects at each point. The total number of objects is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section 3.2 for details.
3.2 Application to Galactic CAGB stars
After confirming the reliability and uncertainties of the CAGB models for LMC-CAGB stars, we extended the analysis to the Milky Way sample (7163 MW-CAGB objects; Table 1), using the 37 CAGB models listed in Table 3 with observed SEDs (see Sections 4 and 4.2 for details).
The top panels of Figure 3 present the WISE–2MASS CMDs for Galactic CAGB stars, plotting M(W3[12]) and M(K[2.2]) against the K[2.2]-W3[12] color. Distances were derived by fitting the observed SEDs with the CAGB models (see Table 3; Section 4). To account for uncertainties in the luminosity of the CAGB models for Galactic CAGB stars, we applied Gaussian Monte Carlo scatter (GS) in distance ($\log_{10}d$) with $\sigma=\pm 0.0798$ dex, consistent with the dispersion measured for LMC-CAGB stars (see the right panel of Figure 1). CAGB stars in the LMC and the Milky Way exhibit very similar distributions on the CMDs (see Figures 2 and 3) and show good overall agreement with the CAGB models, displaying a comparable degree of spread.
@@ 3.2 Application to Galactic CAGB stars
The bottom panels of Figure 3 show the W3[12] band luminosity functions for CAGB stars in both galaxies. They exhibit similar overall shapes, each resembling a combination of two roughly Gaussian components: a main one with M(W3[12]) between 8-8 and 10.5-10.5 mag, and a secondary one comprising luminous sources with M(W3[12]) between 10.5-10.5 and 14-14 mag. Compared with the Milky Way, the LMC contains fewer of these luminous sources (M(W3[12]) <10.5<-10.5 mag), which correspond to “extreme carbon stars” representing the late stage of the CAGB phase (see Gruendl et al. 2008). Among Galactic CAGB stars, MW-CAGB_NI objects, which are generally fainter than MW-CAGB_IC objects (see Section 2.2), are rare at this bright end of the luminosity function.
3.3 Limitations of the Theoretical Models
The dust shell models adopted in this study do not account for gas-phase radiative processes (see Section 3). However, CAGB stars exhibit a variety of gas-phase emission and absorption features in the NIR and MIR bands arising from circumstellar molecules such as C2, C2H2, HCN, and CN (e.g., Le Bertre 2005). Consequently, the discrepancies between the models and observations are expected to be more pronounced at wavelengths where gas-phase radiation is significant.
@@ 3.3 Limitations of the Theoretical Models
4 SEDs and Distances of CAGB stars
Suh (2024) compared observed SEDs with theoretical models for a subset of Galactic CAGB stars. In this work, we extend that analysis to the full samples of CAGB stars in both the LMC and the Milky Way (see Table 1), using the radiative transfer models described in Section 3.
@@ 4 SEDs and Distances of CAGB stars
The observed SEDs of LMC-AGB stars were fitted using a least-squares optimization method, resulting in the 37 CAGB models (Table 3; see Section 3.1 and Figure 1) that best reproduced the observed SED shapes and minimized the differences between the scaling and true distances of LMC-CAGB stars. These LMC-validated CAGB models were then applied to the full sample of Galactic CAGB stars.
@@
4.1 SEDs of LMC-CAGB stars
@@
Figure 4: Observed and model SEDs for CAGB stars. The upper four panels show LMC stars, and the lower four show Milky Way stars. Each panel indicates the $\tau_{10}$ value and the corresponding scaling distance (bottom right). A suffix “s2” on $\tau_{10}$ denotes AMC+SiC dust opacity (20 % SiC). See Section 4 for details.
Figure 5: Observed and model SEDs for two LMC-CAGB stars with thick dust envelopes. The object in the right panel is likely an extrinsic carbon star (see Section 4.1).
Figure 6: Observed and model SEDs for Galactic CAGB stars with very thick dust envelopes ($\tau_{10}>1.5$). Such heavily obscured objects are not seen among LMC-CAGB stars (see Section 4.2).
4.2 SEDs of Galactic CAGB stars

### v4

# Properties of Carbon-rich Asymptotic Giant Branch Stars in the LMC and the Milky Way

@@ Abstract
We present a comparative study of carbon-rich asymptotic giant branch (CAGB) stars in the Large Magellanic Cloud (LMC; 7347 stars) and the Milky Way (7163 stars) using infrared color–magnitude diagrams, spectral energy distributions (SEDs), two-color diagrams, and variability data. Observed SEDs are compared with theoretical models to characterize the central stars and their circumstellar dust envelopes and to estimate distances. For the LMC, a set of best-fitting CAGB models is derived by fitting observed SEDs with radiative transfer models, utilizing the galaxy’s well-established distance. For Galactic CAGB stars, where Gaia DR3 parallaxes are uncertain, we estimate distances by fitting observed SEDs with the CAGB models validated against LMC stars, and for Mira variables, from the period–magnitude relation calibrated with LMC Miras. A comparison of these approaches demonstrates that the SED-based distances are both reliable and practical for a large sample of Galactic CAGB stars. We find that CAGB stars in both galaxies show broadly similar infrared properties, although the LMC sample lacks stars with extremely thick dust envelopes.
@@ I Introduction
LMC-CAGB_SAGE-Sb Sloan et al. (2016); Jones et al. (2017) - 151
LMC-CAGB_KDMK Kontizas et al. (2001) 7760c 24+15d
LMC-CAGB combinede - 7347
@@
MW-CAGB combinede - 7163
aSee Section II.1. bSubgroup of LMC-CAGB_SAGE; objects identified from the Spitzer IRS. c3933 objects are duplicates of LMC-CAGB_SAGE objects. A large fraction of this sample could be extrinsic carbon stars. dMiras from OGLE-III and candidates for new Miras identified from WISE data (Suh 2025). eIt denotes the merged set of all unique objects in the catalogs listed above.
@@
We analyze the properties of CAGB stars in the LMC and the Milky Way. Section II presents catalogs of known CAGB stars in both galaxies. Section III describes radiative transfer models for circumstellar dust shells and the derivation of model parameters, constrained by LMC CAGB stars. In Section IV, observed spectral energy distributions (SEDs) are compared with model predictions to characterize the central stars and their dust envelopes and to estimate theoretical distances. Section V compares observed infrared two-color diagrams (2CDs) and color–magnitude diagrams (CMDs) with theoretical models. Section VI examines the period–magnitude relation (PMR) of Mira-type CAGB stars. Section VII summarizes the main results.
II Sample Stars
@@ II.1 CAGB stars in the LMC
II.2 CAGB stars in the Milky Way
Suh (2024) compiled a catalog of CAGB stars in the Milky Way, classified into two groups: those with IRAS counterparts (MW-CAGB_IC), identified from the IRAS source catalog and generally corresponding to brighter or more isolated stars, and those without IRAS counterparts (MW-CAGB_NI), identified from the AllWISE or Gaia DR3 catalogs and typically representing fainter stars or objects in crowded regions. Most MW-CAGB_IC sources also have entries in the AllWISE or Gaia DR3 catalogs. In this study, we adopt the catalog of Galactic CAGB stars from Suh (2024) (MW-CAGB objects; see Table 1). The MW-CAGB_IC-S subset denotes those MW-CAGB_IC objects for which infrared spectral data are available (see Section IV).
@@ II.2 CAGB stars in the Milky Way
II.3 Observational Data
Table 2 presents the visual and IR bands used in this study for constructing SEDs, CMDs and 2CDs. The table includes the reference wavelength ($\lambda_{ref}$) and zero-magnitude flux (ZMF) for each band, which are essential for deriving theoretical models (see Section III) and building SEDs from observational data. For a comprehensive discussion of IR photometric data and ZMF values, refer to Section 2.1 of Suh (2021).
@@ II.3 Observational Data
II.4 Distances and extinction for Galactic CAGB stars
@@
Gaia detectors operating in the visual bands cannot reliably detect CAGB stars with thick dust envelopes, often resulting in missing data for these objects. In addition, the Gaia counterparts provided by SIMBAD may be inaccurate for some objects, as it is often difficult to identify reliable counterparts using available photometric data and theoretical models, especially in crowded regions. Furthermore, CAGB stars frequently exhibit non-symmetric motions in their outer envelopes (e.g., Kim & Taam 2012), which can lead to inaccurate distance measurements in Gaia DR3. Therefore, distances derived from comparative analyses between observed and theoretical model SEDs may provide alternative or more reliable estimates (see Section IV).
@@
III Theoretical Dust Shell Models
To calculate theoretical model SEDs for CAGB stars, we use radiative transfer models for spherically symmetric dust shells surrounding central stars. The radiative transfer code RADMC-3D (https://github.com/dullemond/radmc3d-2.0), following the approaches of Suh & Kwon (2013) and Suh (2024), is applied. Detailed descriptions of the theoretical models and their limitations can be found in Suh (2020).
@@ III Theoretical Dust Shell Models
aSee Section III for details. For all models, $T_{c}$ = 1000 K. bSeven representative CAGB models were used to produce CMDs and 2CDs. cAMC: pure AMC; AMC+SIC: a simple mixture of AMC and SiC (20% by mass) dStellar luminosity in $10^{3}L_{\odot}$. Based on this luminosity, the scaled logarithmic distances can be sampled using a Gaussian Monte Carlo spread (GS) in $\log_{10}d$ ($\sigma=0.0798$ dex) derived from LMC-CAGB stars (see Figure 1), corresponding to a total luminosity variation of approximately $2.1\times$ from minimum to maximum (see Section III.1).
III.1 Modeling LMC-CAGB stars
@@
For the full sample of 7347 LMC-CAGB stars (Table 1), we tested an extensive grid of radiative transfer models covering a broad range of parameter combinations to identify the most suitable set of CAGB models. The observed SEDs were fitted using a least-squares optimization method (see Sections IV and IV.1).
Figure 1: The top panel shows CAGB model SEDs for seven representative models (see Table 3). For LMC-CAGB stars, the figure shows the distance distribution (middle) and the relations between dust optical depth and luminosity (bottom) derived from CAGB model SEDs (Table 3). See Section III.1 for details.
Figure 2: For LMC-CAGB stars, the top panels show the relations between optical depth and luminosity derived using two different methods based on CAGB models. In the top panels, the areas of the filled circular markers are proportional to the square root of the number of objects at each point. The bottom panels present IR CMDs for LMC-CAGB stars, with the brown arrow indicating an object likely to be an extrinsic carbon star (see Section IV.1). For CAGB models (AMC, $T_{c}$ = 1000 K): $\tau_{10}$ = 0.0001, 0.01, 0.1, 0.5, 1, 2, and 4 from left to right (see Table 3). The total number of objects in each subgroup is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section III.1 for details.
@@
The best-fitting CAGB model (from Table 3) and the corresponding scaling distance were derived for each LMC-CAGB star. Reliable results were obtained for 7205 of the 7347 LMC-CAGB stars (see Section IV). The bottom panel of Figure 1 shows the logarithmic distribution of the derived distances for the 7205 LMC-CAGB stars. The median distance (49.6 kpc) agrees well with the established value for the LMC, and the small dispersion in $\log_{10}d$ (0.0798 dex) demonstrates the consistency and reliability of the CAGB models (Table 3).
@@
The dispersion in $\log_{10}d$ (0.0798 dex; corresponding to roughly 20 % in distance; bottom panel of Figure 1) for LMC-CAGB stars likely arises from three factors: (1) positional variations within the LMC, whose radius is about 10 % of its mean distance, implying a $\sim$5 % dispersion in distance; (2) the fact that a given $\tau_{10}$ value can correspond to multiple luminosities, as stars with different initial masses may exhibit different luminosities; and (3) intrinsic luminosity variations caused by the pulsations of CAGB stars. The observed dispersion in $\log_{10}d$ can therefore be regarded as representing the overall uncertainty in luminosity, since the contribution of the first factor is relatively minor. The measured dispersion in $\log_{10}d$ corresponds to a luminosity dispersion of 0.160 dex for the CAGB models (see Table 3).
Figure 3: The top panels show IR CMDs for Galactic CAGB stars. For Galactic CAGB stars, distances were derived from CAGB model SEDs, with Gaussian Monte Carlo scatter (GS) in $\log_{10}d$ ($\sigma$$\pm$0.0798 dex; obtained from LMC-CAGB stars; see Figure 1). For CAGB stars in the LMC and the Milky Way, the middle panels show the relations between optical depth and luminosity, and the bottom panels show the luminosity functions. In the middle panels, the areas of the filled circular markers are proportional to the square root of the number of objects at each point. The total number of objects is given, with the values in parentheses denoting the number of plotted sources for which data are available. See Section III.2 for details.
III.2 Application to Galactic CAGB stars
After confirming the reliability and uncertainties of the CAGB models for LMC-CAGB stars, we extended the analysis to the Milky Way sample (7163 MW-CAGB objects; Table 1), using the 37 CAGB models listed in Table 3 with observed SEDs (see Sections IV and IV.2 for details).
The top panels of Figure 3 present the WISE–2MASS CMDs for Galactic CAGB stars, plotting M(W3[12]) and M(K[2.2]) against the K[2.2]-W3[12] color. Distances were derived by fitting the observed SEDs with the CAGB models (see Table 3; Section IV). To account for uncertainties in the luminosity of the CAGB models for Galactic CAGB stars, we applied Gaussian Monte Carlo scatter (GS) in distance ($\log_{10}d$) with $\sigma=\pm 0.0798$ dex, consistent with the dispersion measured for LMC-CAGB stars (see the right panel of Figure 1). CAGB stars in the LMC and the Milky Way exhibit very similar distributions on the CMDs (see Figures 2 and 3) and show good overall agreement with the CAGB models, displaying a comparable degree of spread.
@@ III.2 Application to Galactic CAGB stars
The bottom panels of Figure 3 show the W3[12] band luminosity functions for CAGB stars in both galaxies. They exhibit similar overall shapes, each resembling a combination of two roughly Gaussian components: a main one with M(W3[12]) between 8-8 and 10.5-10.5 mag, and a secondary one comprising luminous sources with M(W3[12]) between 10.5-10.5 and 14-14 mag. Compared with the Milky Way, the LMC contains fewer of these luminous sources (M(W3[12]) <10.5<-10.5 mag), which correspond to “extreme carbon stars” representing the late stage of the CAGB phase (see Gruendl et al. 2008). Among Galactic CAGB stars, MW-CAGB_NI objects, which are generally fainter than MW-CAGB_IC objects (see Section II.2), are rare at this bright end of the luminosity function.
III.3 Limitations of the Theoretical Models
The dust shell models adopted in this study do not account for gas-phase radiative processes (see Section III). However, CAGB stars exhibit a variety of gas-phase emission and absorption features in the NIR and MIR bands arising from circumstellar molecules such as C2, C2H2, HCN, and CN (e.g., Le Bertre 2005). Consequently, the discrepancies between the models and observations are expected to be more pronounced at wavelengths where gas-phase radiation is significant.
@@ III.3 Limitations of the Theoretical Models
IV SEDs and Distances of CAGB stars
Suh (2024) compared observed SEDs with theoretical models for a subset of Galactic CAGB stars. In this work, we extend that analysis to the full samples of CAGB stars in both the LMC and the Milky Way (see Table 1), using the radiative transfer models described in Section III.
@@ IV SEDs and Distances of CAGB stars
The observed SEDs of LMC-AGB stars were fitted using a least-squares optimization method, resulting in the 37 CAGB models (Table 3; see Section III.1 and Figure 1) that best reproduced the observed SED shapes and minimized the differences between the scaling and true distances of LMC-CAGB stars. These LMC-validated CAGB models were then applied to the full sample of Galactic CAGB stars.
@@
IV.1 SEDs of LMC-CAGB stars
@@
Figure 4: Observed and model SEDs for CAGB stars. The upper four panels show LMC stars, and the lower four show Milky Way stars. Each panel indicates the $\tau_{10}$ value and the corresponding scaling distance (bottom right). A suffix “s2” on $\tau_{10}$ denotes AMC+SiC dust opacity (20 % SiC). See Section IV for details.
Figure 5: Observed and model SEDs for two LMC-CAGB stars with thick dust envelopes. The object in the right panel is likely an extrinsic carbon star (see Section IV.1).
Figure 6: Observed and model SEDs for Galactic CAGB stars with very thick dust envelopes ($\tau_{10}>1.5$). Such heavily obscured objects are not seen among LMC-CAGB stars (see Section IV.2).
IV.2 SEDs of Galactic CAGB stars